Some lenders offer loans with biweekly payments rather than monthly payments. Investigate the effect of this on home loans by finding the payment and total interest on a thirty-year simple interest amortized loan of $150,000 at 13% interest. (Round your answers to the nearest cent.)

(a) if payments are made monthly

payment will be    $_____
total interest will be $_____


(b) if payments are made biweekly

payment will be    $_____
total interest will be $_____
Answer #1

Amortization payment formula,

p = payment

P = Loan Amount =

n = Number of payment per year

r = Annual interest rate

R =r/n = Periodic Interest rate

t = years

N = Total No. of Payments = n*t

P = 150,000, r=13%, t = 3

a) if payments are made monthly?

n = 12, N = 12*3 = 36, R= 13/12 %

Using amortization payment formula

p = $5054.09

Total payment over three year = 36*p = $181,947.24

Total Interest = 181,947.24 - 150,000 = $31,947.24

Payment = $5054.09

Total Interest = $31,947.24

b) if payments are made biweekly?

n = 52/2 = 26, N = 26*3 = 78, R= 13/26 %

Using amortization payment formula

p = $2327.13

Total payment over three year = 78*p = $181,516.14

Total Interest = 181516.14 - 150,000 = $31,516.14

Payment = $2327.13

Total Interest = $31,516.14